# A general comparison principle for the pluripotential complex Monge-Ampère flow

### v2

@@ 1. Introduction
Let dVdV be a Lebesgue measure on $\mathbb{C}^{n}$. Guedj, Lu and Zeriahi [GLZ21-2] proved the existence and uniqueness of the solution to the Cauchy-Dirichlet problem with a measure $\mu$ and a boundary data $h$ satisfying
@@
Organization. In Section 22, we recall the definition of parabolic Monge-Ampère operator in [GLZ21-2]. We also prove several useful lemmas about the operator. In Section 33, we first prove some lemmas about mixed type inequalities and property of supersolutions. Then we give the proof of Theorem 1.2 in Theorem 3.8 by using these lemmas. With this result, we also prove the uniqueness of the weak solution to the Cauchy-Dirichlet problem and study the long-term behavior of the solution.
@@ 2. Parabolic Monge-Ampère Operator
It follows from Lemma 2.2 that each side of (2.7) is a Borel measurable function on $(0,T)$. By integrating each side with respect to dtdt for $t\in(0,T)$,
@@ 3.2. Boundary behavior of supersolutions
for all $t,s\in[t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$] and $z\in\Omega$. It follows from the assumption (c) that there exists $C_{2}>0$ such that $u(t,z)-C_{2}t^{2}$ and $v(t,z)-C_{2}t^{2}$ are concave in $t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$] for all $z\in\Omega$. Let us define $C_{3}:=\max\{C_{1},C_{2}\}$, so that we have
@@
for all $t,s\in[t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$] and $z\in\Omega$, and
(3.2) $u(t,z)-C_{3}t^{2},v(t,z)-C_{3}t^{2}\text{ are concave in }[t_{1}-\tilde{\varepsilon},t_{1}+\tilde{\varepsilon}$]
@@
Let us denote $\delta:=\min\{\frac{\tilde{\varepsilon}}{2},\frac{\varepsilon}{6C_{3}}\}$ and fix $t\in(t_{1}-\delta,t_{1}+\delta)$. Since $\mu(E_{t-\delta})=0$, we have
@@
[…]
@@
for some $B>0$ in $\Omega_{S}$. We will remove this assumption at the end of the proof. For $s>0$ near 11 we set, for $(t,z)\in\Omega_{S}$,
@@
and apply Lemma 3.6. Note that (i) follows from Lemma $3.$ To prove (ii), it suffices to show that for each open $U\Subset\Omega$, there exists $A_{0}>0$ such that

### v3

@@ 1. Introduction
Let $dV$ be a Lebesgue measure on $\mathbb{C}^{n}$. Guedj, Lu and Zeriahi [GLZ21-2] proved the existence and uniqueness of the solution to the Cauchy-Dirichlet problem with a measure $\mu$ and a boundary data $h$ satisfying
@@
Organization. In Section $2$, we recall the definition of parabolic Monge-Ampère operator in [GLZ21-2]. We also prove several useful lemmas about the operator. In Section $3$, we first prove some lemmas about mixed type inequalities and property of supersolutions. Then we give the proof of Theorem 1.2 in Theorem 3.8 by using these lemmas. With this result, we also prove the uniqueness of the weak solution to the Cauchy-Dirichlet problem and study the long-term behavior of the solution.
@@ 2. Parabolic Monge-Ampère Operator
It follows from Lemma 2.2 that each side of (2.7) is a Borel measurable function on $(0,T)$. By integrating each side with respect to $dt$ for $t\in(0,T)$,
@@ 3.3. Main result
for all $t,s\in\left[t_{1}-\frac{\tilde{\varepsilon}}{2},t_{1}+\frac{\tilde{\varepsilon}}{2}\right]$ and $z\in\Omega$. It follows from the assumption (c) that there exists $C_{2}>0$ such that $u(t,z)-C_{2}t^{2}$ and $v(t,z)-C_{2}t^{2}$ are concave in $\left[t_{1}-\frac{\tilde{\varepsilon}}{2},t_{1}+\frac{\tilde{\varepsilon}}{2}\right]$ for all $z\in\Omega$. Let us define $C_{3}:=\max\{C_{1},C_{2}\}$, so that we have
@@
for all $t,s\in\left[t_{1}-\frac{\tilde{\varepsilon}}{2},t_{1}+\frac{\tilde{\varepsilon}}{2}\right]$ and $z\in\Omega$, and
(3.2) $u(t,z)-C_{3}t^{2},v(t,z)-C_{3}t^{2}\text{ are concave in }\left[t_{1}-\frac{\tilde{\varepsilon}}{2},t_{1}+\frac{\tilde{\varepsilon}}{2}\right]$
@@
Let us denote $\delta:=\min\{\frac{\tilde{\varepsilon}}{4},\frac{\varepsilon}{6C_{3}}\}$ and fix $t\in(t_{1}-\delta,t_{1}+\delta)$. Since $\mu(E_{t-\delta})=0$, we have
@@
Note that $t$ and $t-\delta$ are in $\left(t_{1}-\frac{\tilde{\varepsilon}}{2},t_{1}+\frac{\tilde{\varepsilon}}{2}\right)$. By combining (3.1) and (3.4), we get
@@
for some $B>0$ in $\Omega_{S}$. We will remove this assumption at the end of the proof. For $s>0$ near $1$ we set, for $(t,z)\in\Omega_{S}$,
@@
and apply Lemma 3.6. Note that (i) follows from Lemma 3.3. To prove (ii), it suffices to show that for each open $U\Subset\Omega$, there exists $A_{0}>0$ such that